### v1

@@ Abstract
This paper introduces MultiBooth, a novel and efficient technique for multi-concept customization in image generation from text. Despite the significant advancements in customized generation methods, particularly with the success of diffusion models, existing methods often struggle with multi-concept scenarios due to low concept fidelity and high inference cost. MultiBooth addresses these issues by dividing the multi-concept generation process into two phases: a single-concept learning phase and a multi-concept integration phase. During the single-concept learning phase, we employ a multi-modal image encoder and an efficient concept encoding technique to learn a concise and discriminative representation for each concept. In the multi-concept integration phase, we use bounding boxes to define the generation area for each concept within the cross-attention map. This method enables the creation of individual concepts within their specified regions, thereby facilitating the formation of multi-concept images. This strategy not only improves concept fidelity but also reduces additional inference cost. MultiBooth surpasses various baselines in both qualitative and quantitative evaluations, showcasing its superior performance and computational efficiency.
Keywords:
Text-to-image generation personalized image generation multi-concept customization
Figure 1: MultiBooth can learn individual customization concepts through a few examples and then combine these learned concepts to create multi-concept images based on text prompts. The results indicate that our MultiBooth can effectively preserve high image fidelity and text alignment when encountering complex multi-concept generation demands, including (a) restylization, (b) different spatial relationships, and (c) recontextualization.
@@ 1 Introduction
The advent of diffusion models has ignited a new wave in the text-to-image (T2I) task, leading to the proposal of numerous models [23, 17, 34, 5, 31, 4, 9, 13, 28]. Despite the broad capabilities of these models, users often desire to generate specific concepts such as beloved pets or personal items. These personal concepts are not captured during the training of large-scale T2I models due to their subjective nature, emphasizing the need for customized generation [24, 1, 29, 8, 10, 2, 14, 16, 15]. Customized generation aims to create new variations of a given concept, including different contexts (e.g., beaches, forests) and styles (e.g., painting), based on just a few user-provided images (typically fewer than 5).
Recent customized generation methods either learn a concise token representation for each subject [17] or adopt an efficient fine-tuning strategy to adapt the T2I model specifically for the subject [24]. While these methods have achieved impressive results, they primarily focus on single-concept customization and struggle when users want to generate customized images for multiple subjects (see Fig. 2). This motivates the study of multi-concept customization (MCC).
Existing methods[8] for MCC commonly employ joint training approaches. However, this strategy often leads to feature confusion, as illustrated in the third column of Fig. 2. Furthermore, these methods require training distinct models for each combination of subjects and are hard to scale up as the number of subjects grows. An alternative method [10] addresses MCC by adjusting attention maps with residual token embeddings during inference. While this approach shows promise, it incurs a notable inference cost. Furthermore, the method encounters difficulties in attaining high fidelity due to the restricted learning capacity of a single residual embedding.
To address the aforementioned issues, we introduce MultiBooth, a two-phase MCC solution that accurately and efficiently generates customized multi-concept images based on user demand, as demonstrated in the Fig. 2. MultiBooth includes a discriminative single-concept learning phase and a plug-and-play multi-concept integration phase. In the former phase, we learn each concept separately, resulting in a single-concept module for every concept. In the latter phase, we effectively combine these single-concept modules to generate multi-concept images without any extra training.
@@
In the plug-and-play multi-concept integration phase, we further propose a regional customization module to guide the inference process, allowing the correct combination of different single-concept modules for multi-concept image generation. Specifically, we divide the attention map into different regions within the cross-attention layers of the U-Net, and each region’s attention value is guided by the corresponding single-concept module and prompt. Through the proposed regional customization module, we can generate multi-concept images via any combination of single-concept modules while bringing minimal cost during inference.
Figure 2: Results of single-concept methods (Text-Inversion [1] and DreamBooth [24]) and multi-concept methods (Custom Diffusion [8], Cone2 [10], and our MultiBooth). Traditional single-concept methods are inadequate when it comes to accommodating multi-concept customization, whereas existing multi-concept approaches often struggle with maintaining high fidelity and prompt alignment.
@@ 2 Related Work
2.1 Text to Image Generation
Text-to-image generation is an extensively researched problem, with numerous early studies focusing on Generative Adversarial Networks (GANs)[3]. Noteworthy examples include AttnGAN[30], StackGAN[32], StackGAN++[33], Mirrorgan[20]. While these methods have proven to be effective on specific datasets such as faces and landscapes, they have limited generalization ability to larger-scale datasets. Moreover, training GANs poses instability issues and is susceptible to mode collapse. With the advancements of diffusion models[6], more and more methods have started exploring text-to-image generation based on diffusion models. By training on large-scale text-image datasets such as LAION[26], these methods have achieved superior performance. Notable recent works include DALLE2[22], Imagen[25], GLIDE[18], and Stable Diffusion[23], which are capable of generating images based on open-vocabulary text prompts.
2.2 Customized Text to Image Generation
The goal of customized text-to-image generation is to acquire knowledge of a novel concept from a limited set of examples and subsequently generate images of these concepts in diverse scenarios based on text prompts. By leveraging the aforementioned diffusion-based methodologies, it becomes possible to employ the comprehensive text-image prior to customizing the text-to-image process.
Textual Inversion[1] achieves customization by creating a new embedding in the tokenizer and associating all the details of the newly introduced concept to this embedding. DreamBooth[24] binds the new concept to a rare token followed by a class noun. This process is achieved by finetuning the entire diffusion model. Additionally, DreamBooth addresses the issue of language drift through a prior preservation loss. ELITE[29] utilizes multi-layer embeddings from CLIP image encoder[21] and employs a mapping network to acquire customized embeddings. Similar approaches are adopted in InstantBooth[27] and E4T[2]. Such encoder-based methods require less training time and yield better image generation results compared to directly optimizing the embedding in Textual Inversion[1]. Custom Diffusion[8] links a novel concept to a rare token by adjusting specific parameters of the diffusion model. To alleviate overfitting, Custom Diffusion also incorporates a regularization dataset. Through joint training, Custom Diffusion explores the problem of multi-concept customization for the first time. Cones2[10] learns new concepts by adding a residual embedding on top of the base embedding and generates multi-concept images through attention map manipulation.
In this work, we utilize a multi-modal model and LoRA to discriminatively and concisely encode every single concept. Then, we introduce the regional customization module to efficiently and accurately produce multi-concept images.
@@ 3 Method
Given a series of images $\mathcal{S}=\{X_{s}\}^{S}_{s=1}$ that represent $S$ concepts of interest, where $\{X_{s}\}=\{x_{i}\}^{M}_{i=1}$ denotes the $M$ images belonging to the concept $s$ which is usually very small (e.g., $M<=5$), the goal of multi-concept customization (MCC) is to generate images that include any number of concepts from $\mathcal{S}$ in various styles, contexts, layout relationship as specified by given text prompts. MCC poses significant challenges for two primary reasons. Firstly, learning a concept with a limited number of images is inherently difficult. Secondly, generating multiple concepts simultaneously and coherently within the same image while faithfully adhering to the provided text is even harder.
To tackle these challenges, Custom Diffusion[8] simultaneously finetunes the model with all subjects of interest. This approach can lead to issues like feature confusion and inaccurate attribution (e.g., attributing dog features to a cat). For Cone2[10], the process requires repetitive strengthening and weakening of the same concept features during the inference phase, which can result in a significant decrease in fidelity. To avoid these problems, we propose a method called MultiBooth. Our MultiBooth initially performs high-fidelity learning of a single concept. We employ a multi-modal encoder and the adaptive concept normalization strategy to obtain text-aligned representative customized embeddings. Additionally, the efficient concept encoding technique is employed to further improve the fidelity of single-concept learning. To generate multi-concept images, we employ the regional customization module. This module serves as a guide for multiple single-concept modules and utilizes bounding boxes to indicate the positions of each generated concept.
@@ 3.1 Preliminaries
In this paper, the foundational model utilized for text-to-image generation is Stable Diffusion [23]. It takes a text prompt $P$ as input and generates the corresponding image $x$. Stable diffusion[23] consists of three main components: an autoencoder$(\mathcal{E}(\cdot),\mathcal{D}(\cdot))$, a CLIP text encoder $\tau_{\theta}(\cdot)$ and a U-Net $\epsilon_{\theta}(\cdot)$. Typically, it is trained with the guidance of the following reconstruction loss:
@@
Given $M$ images $\{X_{s}\}=\{x_{i}\}_{i=1}^{M}$ of a certain concept $s$, previous works[1, 24, 8] associate a unique placeholder string $S^{*}$ with concept $s$ through a specific prompt $P_{s}$ like “a photo of a $S^{*}$ dog”, with the following finetuning objective:
@@
The result of minimizing Eq. 2 is to encourage the U-Net $\epsilon_{\theta}(\cdot)$ to accurately reconstruct the images of the concept $s$, effectively binding the placeholder string $S^{*}$ to the concept $s$.
Figure 3: Overall Pipeline of MultiBooth. (a) During the single-concept learning phase, a multi-modal encoder and LoRA parameters are trained to encode every single concept. (b) During the multi-concept integration phase, the customized embeddings $S^{*}$ and $V^{*}$ are converted into text embeddings, which are then combined with the corresponding LoRA parameters to form single-concept modules. These single-concept modules, along with the bounding boxes, are intended to serve as input for the regional customization module.
@@ 3.2 Single-Concept Learning
3.2.1 Multi-modal Concept Extraction.
Existing methods[2, 29] mainly utilize a single image encoder to encode the concepts of interest. However, this approach may also encode irrelevant information, such as unrelated objects in the images. To remedy this, we employ a multi-modal encoder that takes as input both the images and the subject name (e.g., “dog”) to learn a concise and discriminative representation for each concept. Inspired by MiniGPT4 [35] and BLIP-Diffusion [9], we utilize the QFormer, a light-weighted multi-modal encoder, to generate customized concept embeddings. The QFormer encoder $E$ has three types of inputs: visual embeddings $\xi$ of an image, text description $l$, and learnable query tokens $W=[w_{1},\cdots,w_{K}]$ where $K$ is the number of query tokens. The outputs of QFormer are tokens $O=[o_{1},\cdots,o_{K}]$ with the same dimensions as the input query tokens.
As shown in Fig. 3, given an image $x_{i}\in X_{s}$, we employ a frozen CLIP[21] image encoder to extract the visual embeddings $\xi$ of the image. Subsequently, we set the input text $l$ as the subject name for the image and input it into the encoder $E$. The learnable query tokens $W$ interact with the text description $l$ through a self-attention layer and with the visual embedding $\xi$ through a cross-attention layer, resulting in text-image aligned output tokens $O=E(\xi,l,W)$. Finally, we obtain the initial customized embedding $v_{i}$ by taking the average of these tokens:
$v_{i}=\frac{1}{K}\cdot\sum_{i=1}^{K}{o_{i}}$ (3)
where $v_{i}\in\mathbb{R}^{d}$, $d$ is the dimension of our customized embedding. Following [1, 24, 8], a placeholder string $S^{*}$ is introduced to represent the newly acquired concept, with $v_{i}$ representing its word embedding. Through this placeholder string $S^{*}$, we can easily activate the customized word embedding $v_{i}$ to reconstruct the input concept image $x_{i}$ with prompts like “a photo of a $S^{*}$ dog”.
Table 1: Quantization results of the L2 norm of each word embedding in the prompt. The customized embeddings of Textual Inversion and Ours w/o ACN have significantly larger L2 norm compared to other word embeddings in the prompt.
@@
3.2.2 Adaptive Concept Normalization
During the reconstruction with the prompt mentioned above, we have observed a domain gap between our customized embedding $v_{i}$ and other word embeddings in the prompt. As shown in Tab. 1, the L2 norm of our customized embedding is considerably larger than that of other word embeddings in the prompt. Notably, these word embeddings, belonging to the same order of magnitude, are predefined within the embedding space of the CLIP text encoder $\tau_{\theta}(\cdot)$. This significant difference in quantity weakens the model’s ability for multi-concept generation. To remedy this, we further apply the Adaptive Concept Normalization (ACN) strategy to the customized embedding $v_{i}$, adjusting its L2 norm to obtain the final customized embedding $\hat{v_{i}}$
The adaptive concept normalization strategy consists of two steps. The first step is L2 normalization, adjusting the L2 norm of the customized embedding $v_{i}$ to $1$. The second step is adaptive scaling, which brings the L2 norm of $v_{i}$ to a comparable magnitude as other word embeddings in the prompt. Specifically, let $c_{l}\in\mathbb{R}^{d}$ represent the word embedding corresponding to the subject name of $v_{i}$ (e.g., the word embedding of “dog”), where $d$ is the dimension of embeddings. The adaptive concept normalization can be expressed as:
$\hat{v_{i}}=v_{i}\cdot\frac{\lVert c_{l}\rVert_{2}}{\lVert v_{i}\rVert_{2}}$ (4)
where $\hat{v_{i}}\in\mathbb{R}^{d}$ has the same dimensions as $v_{i}$. As shown in Tab. 1, this operation effectively addresses the problem of domain gap in the embedding space.
@@ 3.2.2 Adaptive Concept Normalization
Figure 4: Regional Customization Module. To generate individual concepts within their specified regions, we initially divide the image feature into several regions via bounding boxes to acquire the query $Q$ for each concept. Subsequently, we combine the single-concept module with $W_{k}$ and $W_{v}$ to derive the corresponding key $K$ and value $V$. After that, we perform the attention operation on the obtained $Q$, $K$, and $V$ to get a part of the final output. The above procedure is applied to each concept simultaneously, forming the final output attention map.
3.2.3 Efficient Concept Encoding
@@
Given the modified customized embedding $\hat{v_{i}}$, we can easily reconstruct the input image $x_{i}$ through the text prompt $P_{s}$ containing the placeholder string $S^{*}$ (e.g., “a photo of a $S^{*}$ dog”). To further improve the reconstruction fidelity during single-concept learning and avoid language drift caused by finetuning the U-Net, we incorporate the LoRA technique[7] for efficient concept encoding. Specifically, we incorporate a low-rank decomposition to the key and value weight matrices of attention layers within the U-Net $\epsilon_{\theta}(\cdot)$. Each pre-trained weight matrix $W_{init}\in\mathbb{R}^{d\times k}$ is utilized in the forward computation as follows:
$h=W_{init}x+\Delta Wx=W_{init}x+BAx$ (5)
@@
where $A\in\mathbb{R}^{r\times k},\ B\in\mathbb{R}^{d\times r}$ are trainable parameters of efficient concept encoding, and the rank $r\ll\min(d,k)$. During training, the pre-trained weight matrix $W_{init}$ stays constant without receiving gradient updates.
The whole single-concept learning framework can be trained by optimizing Eq. 2 with a regularization term, as shown in the following equation:
@@
$\mathcal{L}=\mathbb{E}_{z=\mathcal{E}(x),x\sim X_{s},\epsilon,t,P_{s}}\left[\lVert\epsilon-\epsilon_{\theta}\left(z_{t},t,\tau_{\theta}\left(P_{s}\right)\right)\rVert_{2}^{2}\right]+\lambda\lVert\hat{v_{i}}\rVert_{2}^{2}$ (6)
where $\lambda$ denotes a balancing hyperparameter and is consistently set to 0.01 across all experiments. During training, we randomly select text prompts $P_{s}$ from the CLIP ImageNet templates [21] following the Textual Inversion [1]. The complete templates can be found in the $Suppl$. Through the above methods, the detailed information of a new concept can be extracted and stored in a customized embedding and corresponding LoRA parameters, which can be called a single-concept module. The storage requirement of a concept is less than 7MB, which is significantly advantageous compared to 3.3GB in [24] and 72MB in Custom Diffusion[8].

### v2

@@ Abstract
This paper introduces MultiBooth, a method that generates images from texts containing various concepts from users. Despite diffusion models bringing significant advancements for customized text-to-image generation, existing methods often struggle with multi-concept scenarios due to low concept fidelity and high inference cost. MultiBooth addresses these issues by dividing the multi-concept generation process into two phases: a single-concept learning phase and a multi-concept integration phase. During the single-concept learning phase, we employ a multi-modal image encoder and an efficient concept encoding technique to learn a concise and discriminative representation for each concept. In the multi-concept integration phase, we use bounding boxes to define the generation area for each concept within the cross-attention map. This method enables the creation of individual concepts within their specified regions, thereby facilitating the formation of multi-concept images. This strategy not only improves concept fidelity but also reduces additional inference cost. MultiBooth surpasses various baselines in both qualitative and quantitative evaluations, showcasing its superior performance and computational efficiency. Project page is on https://multibooth.github.io/.
@@ 1 Introduction
The advent of diffusion models (Ramesh et al. 2022; Saharia et al. 2022; Nichol et al. 2021; He et al. 2023a, 2024c) has ignited a new wave in the text-to-image (T2I) task, leading to numerous novel methods (Hertz et al. 2022; Ye et al. 2023; Gu et al. 2023; Wang et al. 2024d, a, b). Despite their broad capabilities, users often desire to generate specific concepts such as beloved pets or personal items. These personal concepts are not captured during the training of large-scale T2I models due to their subjective nature, emphasizing the need for customized generation (Wei et al. 2023; Gal et al. 2023; Yan et al. 2023; Li et al. 2024; Zhu et al. 2024). Customized generation aims to create new variations of given concepts, including different contexts (e.g., beaches, forests) and styles (e.g., painting), based on just a few user-provided images (typically fewer than 5).
Recent customized generation methods either learn a concise token representation for each subject (Gal et al. 2022) or adopt a fine-tuning strategy to adapt the T2I model specifically for the subject (Ruiz et al. 2023). While these methods have achieved impressive results, they primarily focus on single-concept customization and struggle when users want to generate customized images for multiple subjects. This motivates the study of multi-concept customization (MCC).
Figure 1: MultiBooth can learn individual customization concepts through a few examples and then combine these learned concepts to create multi-concept images based on text prompts. The results indicate that our MultiBooth can effectively preserve high image fidelity and text alignment when encountering complex multi-concept generation demands, including (a) stylization, (b) different spatial relationships, and (c) contextualization.
Existing methods (Kumari et al. 2023) for MCC commonly employ joint training approaches. However, this strategy often leads to feature confusion. Furthermore, these methods require training distinct models for each combination of subjects and are hard to scale up as the number of subjects grows. An alternative method (Liu et al. 2023) addresses MCC by adjusting attention maps with residual token embeddings during inference. While this approach shows promise, it incurs a notable inference cost. Furthermore, the method encounters difficulties in attaining high fidelity due to the restricted learning capacity of a single residual embedding.
To address the aforementioned issues, we introduce MultiBooth, a two-phase MCC solution that accurately and efficiently generates customized multi-concept images based on user demand. MultiBooth includes a discriminative single-concept learning phase and a plug-and-play multi-concept integration phase. In the former phase, we learn each concept separately, resulting in a single-concept module for every concept. In the latter phase, we effectively combine these single-concept modules to generate multi-concept images without any extra training.
@@
In the plug-and-play multi-concept integration phase, we further propose a regional customization module to guide the inference process, allowing the correct combination of different single-concept modules for multi-concept image generation. Specifically, we divide the attention map into different regions within the cross-attention layers of the U-Net, and each region’s attention value is guided by the corresponding single-concept module and prompt. Through the proposed regional customization module, we can generate multi-concept images via any combination of single-concept modules while bringing minimal cost during inference. Fig. 1 shows some examples.
@@
Figure 2: Overall Pipeline of MultiBooth. (a) During the single-concept learning phase, a multi-modal encoder and LoRA parameters are trained to encode every single concept. (b) During the multi-concept integration phase, we first convert $S^{*}$ and $V^{*}$ into text embeddings, which are then combined with the corresponding LoRA to form single-concept modules. These single-concept modules, along with the bounding boxes, are intended to serve as input for the regional customization module.
@@ 2 Related Work
Layout-guided text to image generation. T2I models have benefited numerous new tasks (Ma et al. 2024b, c, 2022, 2023, d; He et al. 2024b, 2023b, 2023c; Fang et al. 2024; Zhong et al. 2024b; Tang et al. 2024, 2023a, 2023b; Chen et al. 2024; Feng et al. 2024; Wang et al. 2024c; Zhong et al. 2024a, c). To achieve finer control that cannot be accomplished using only text prompts, many T2I methods incorporate layout as an additional input to guide the generation process. One branch of these methods (Xie et al. 2023; Phung, Ge, and Huang 2024; Chen, Laina, and Vedaldi 2024; Ma et al. 2024a) involves designing an extra loss function to update the latent variables and guide the sampling process. While these methods can achieve image generation in a single forward pass, their fidelity is inadequate when dealing with complex object interactions or attributes. The other branch of methods (Lian et al. 2023; Bar-Tal et al. 2023; Jiménez 2023) performs denoising separately for each layout and subsequently fuses the results, leading to high computational costs. Different from the aforementioned methods, our method processes all layouts simultaneously, thereby eliminating the need for additional loss functions to guide sampling. Furthermore, our method can effectively handle complex object interactions while maintaining high image fidelity and precise text alignment.
Customized text to image generation. The goal of customized text-to-image generation is to acquire knowledge of a novel concept from a limited set of examples and subsequently generate images of these concepts in diverse scenarios based on text prompts. By leveraging the aforementioned diffusion-based methodologies, it becomes possible to employ the comprehensive text-image prior to customizing the text-to-image process. The first branch of methods (Gal et al. 2022; Chen et al. 2023; Liu et al. 2023) achieves customization by creating a new embedding within the tokenizer and associating all the details of the newly introduced concept to this embedding. The second branch of methods (Wei et al. 2023; Shi et al. 2023; Gal et al. 2023) trains an adapter to generate embeddings. They need strong GPUs and large datasets for training and only support single-concept customization. To adapt to MCC, they need numerous multi-concept images and costly retraining. The third branch of methods (Ruiz et al. 2023; Kumari et al. 2023) binds the new concept to a rare token followed by a class noun. Compared to the previous two branches of methods, they often achieve the best image fidelity. However, this process is achieved by fine-tuning the entire or partial UNet. As a result, they require a larger amount of parameters to store a new concept. In this work, we utilize a multi-modal model and LoRA to discriminatively and concisely encode every single concept. Then, we introduce the regional customization module to efficiently and accurately produce multi-concept images.
@@ 3 Method
Given a series of images $\mathcal{S}=\{X_{s}\}^{S}_{s=1}$ that represent $S$ concepts of interest, where $\{X_{s}\}=\{x_{i}\}^{M}_{i=1}$ denotes the $M$ images belonging to the concept $s$ which is usually very small (e.g., $M<=5$), the goal of multi-concept customization (MCC) is to generate images that include any number of concepts from $\mathcal{S}$ in various styles, contexts, layout relationship as specified by given text prompts.
MCC faces significant challenges for two primary reasons. Firstly, learning a concept with a limited number of images is inherently difficult. Secondly, generating multiple concepts simultaneously and coherently within the same image while faithfully adhering to the provided text is even harder. To address these challenges, our MultiBooth initially performs high-fidelity learning of a single concept. We employ a multi-modal encoder and the adaptive concept normalization strategy to obtain text-aligned representative customized embeddings. Additionally, the efficient concept encoding technique is employed to further improve the fidelity of single-concept learning. To generate multi-concept images, we employ the regional customization module. This module serves as a guide for multiple single-concept modules and utilizes bounding boxes to indicate the positions of each generated concept.
@@ 3.1 Preliminaries
In this paper, the foundational model utilized for text-to-image generation is Stable Diffusion (Rombach et al. 2022). It takes a text prompt $P$ as input and generates the corresponding image $x$. Stable Diffusion (Rombach et al. 2022) consists of three main components: an autoencoder$(\mathcal{E}(\cdot),\mathcal{D}(\cdot))$, a CLIP text encoder $\tau_{\theta}(\cdot)$ and a U-Net $\epsilon_{\theta}(\cdot)$. Typically, it is trained with the guidance of the following reconstruction loss:
@@
Given $M$ images $\{X_{s}\}=\{x_{i}\}_{i=1}^{M}$ of a certain concept $s$, previous works (Gal et al. 2022; Ruiz et al. 2023; Kumari et al. 2023) associate a unique placeholder string $S^{*}$ with concept $s$ through a specific prompt $P_{s}$ like “a photo of a $S^{*}$ dog”, with the following finetuning objective:
@@
Minimizing Eq. 2 can encourage the U-Net $\epsilon_{\theta}(\cdot)$ to accurately reconstruct the images of the concept $s$, effectively binding the placeholder string $S^{*}$ to the concept $s$.
@@ 3.2 Single-Concept Learning
Multi-modal Concept Extraction.
Existing customization methods (Gal et al. 2023; Wei et al. 2023) mainly utilize a single image encoder to encode the whole image into concept embeddings. However, the single image encoder may also encode unrelated objects in the images. To remedy this, we employ a multi-modal encoder that takes as input both the images and the concept name (e.g., “dog”) to generate concise and discriminative customized embeddings.
Inspired by MiniGPT4 (Zhu et al. 2023) and BLIP-Diffusion (Li, Li, and Hoi 2023), we utilize the QFormer, a light-weighted multi-modal encoder, to generate the customized embeddings for each concept. As shown in the left part of Fig. 2, the QFormer encoder $E$ has three types of inputs: visual embeddings $\xi$ of an image, text description $l$ of the concept of interest, and learnable query tokens $W=[w_{1},\cdots,w_{K}]$ where $K$ is the number of query tokens. Given an image $x_{i}\in X_{s}$ of concept $s$, we employ a frozen CLIP (Radford et al. 2021) image encoder to extract the visual embeddings $\xi$ of the image. Subsequently, we set the input text $l$ as the concept name for the image. The learnable query tokens $W$ interact with the text description $l$ through a self-attention layer and with the visual embedding $\xi$ through a cross-attention layer. This interaction results in text-image aligned output tokens $O=E(\xi,l,W)$ with the same dimensions as $W$. Finally, we average these tokens and get initial customized embedding $v_{i}=\frac{1}{K}\cdot\sum_{i=1}^{K}{o_{i}}$.
After obtaining the customized embedding $v_{i}$ of concept $s$, we introduce a placeholder string $S^{*}$ to represent the concept $s$, with $v_{i}$ representing the word embedding of $S^{*}$. Through this placeholder string $S^{*}$, we can easily activate the customized word embedding $v_{i}$ to reconstruct the input concept image $x_{i}$ with prompts like “a photo of a $S^{*}$ dog”.
Table 1: Quantization results of the L2 norm of each word embedding in the prompt.
@@
Figure 3: Regional Customization Module. We initially divide the image feature into several regions via bounding boxes to acquire the query $Q$ for each concept. Subsequently, we combine the single-concept module with $W_{k}$ and $W_{v}$ to derive the corresponding key $K$ and value $V$. After that, we perform the attention operation on the obtained $Q$, $K$, and $V$ to get a partial attention output. The above procedure is applied to each concept simultaneously, forming the final attention output.
Adaptive Concept Normalization.
We have observed a domain gap between our customized embedding $v_{i}$ and other word embeddings in the prompt. As shown in Tab. 1, the L2 norm of our customized embedding is considerably larger than that of other word embeddings in the prompt. Notably, these word embeddings, belonging to the same order of magnitude, are predefined within the embedding space of the CLIP text encoder $\tau_{\theta}(\cdot)$. This significant difference in quantity weakens the model’s ability of multi-concept generation. To remedy this, we further apply the Adaptive Concept Normalization (ACN) strategy to the customized embedding $v_{i}$, adjusting its L2 norm to obtain the final customized embedding $\hat{v_{i}}$.
Our ACN consists of two steps. The first step is L2 normalization, adjusting the L2 norm of the customized embedding $v_{i}$ to $1$. The second step is adaptive scaling, which brings the L2 norm of $v_{i}$ to a comparable magnitude as other word embeddings in the prompt. Specifically, let $c_{l}\in\mathbb{R}^{d}$ represent the word embedding corresponding to the subject name of $v_{i}$ (e.g., the word embedding of “dog”), where $d$ is the dimension of embeddings. The adaptive concept normalization $\hat{v_{i}}=v_{i}\cdot\frac{\lVert c_{l}\rVert_{2}}{\lVert v_{i}\rVert_{2}}$. As shown in Tab. 1, this operation effectively addresses the problem of domain gap in the embedding space.
Efficient Concept Encoding.
@@
To further improve the concept fidelity during single-concept learning and avoid language drift caused by finetuning the U-Net, we incorporate the LoRA technique (Hu et al. 2021; He et al. 2024a) for efficient concept encoding. Specifically, we incorporate a low-rank decomposition to the key and value weight matrices of attention layers within the U-Net $\epsilon_{\theta}(\cdot)$. Each pre-trained weight matrix $W_{init}\in\mathbb{R}^{d\times k}$ of the U-Net $\epsilon_{\theta}(\cdot)$ is utilized in the forward computation as follows:
$h=W_{init}x+\Delta Wx=W_{init}x+BAx,$ (3)
@@
where $A\in\mathbb{R}^{r\times k},\ B\in\mathbb{R}^{d\times r}$ are trainable parameters of efficient concept encoding, and the rank $r\ll\min(d,k)$. During training, the pre-trained weight matrix $W_{init}$ stays constant without receiving gradient updates. We also use a regularization term to lower the L2 norm of $v_{i}$ before ACN. Without this term, the L2 norm of $v_{i}$ can grow large as shown in Tab. 1. Scaling $v_{i}$ with ACN could greatly alter its magnitude, causing information loss. As a result, the whole single-concept learning framework can be trained as follows:
@@
$\mathcal{L}=\mathbb{E}_{z=\mathcal{E}(x),x\sim X_{s},\epsilon,t,P_{s}}\left[\lVert\epsilon-\epsilon_{\theta}\left(z_{t},t,\tau_{\theta}\left(P_{s}\right)\right)\rVert_{2}^{2}\right]+\lambda\lVert{v_{i}}\rVert_{2}^{2},$ (4)
where $\lambda$ denotes a balancing hyperparameter and is consistently set to 0.01 across all experiments.
So far, we can learn a new concept efficiently and store its information in a dedicated single-concept module. This module contains a customized embedding along with the corresponding LoRA parameters. The extra parameter for a new concept is less than 7MB, which is significantly lower compared to 3.3GB in DreamBooth (Ruiz et al. 2023) and 72MB in Custom Diffusion (Kumari et al. 2023). Furthermore, the single-concept module is plug-and-play for multi-concept generation, as users can combine any single-concept module through the Regional Customization Module to perform multi-concept generation.
Figure 4: Qualitative comparisons. Our method outperforms all the compared methods in image fidelity and prompt alignment.